If 15.0 of phosphoric acid completely neutralizes 38.5mL of 0.150 mol/L sodium hydroxide, what is the concentration of the phosphoric acid


1
Expert's answer
2022-05-24T17:36:03-0400

The amount of mole at the neutralization point is equal:

"n_1=n_2"

"c_1V_1=c_2V_2"

"c_1 = \frac{c_2V_2}{V_1} = \frac{0.150 mol/L\times 38.5 ml}{15.0 ml} = 0.385 mol/L"